OpenStudy (anonymous):

How do I expand (2g+2)^8

OpenStudy (shayaan_mustafa):

do you know binomial theorem?

OpenStudy (anonymous):

no :(

OpenStudy (shayaan_mustafa):

hmmm... it is much long and complicated. Do you know what is factorial?

OpenStudy (shayaan_mustafa):

here is the link. and try to understand. i will surely help you. http://www.intmath.com/series-binomial-theorem/4-binomial-theorem.php

OpenStudy (anonymous):

thank you

OpenStudy (shayaan_mustafa):

welcome.

OpenStudy (anonymous):

If I were to expand (3-t)^3 is it 3^3+3(3^2-t)+3(3-t^2)-t^2

OpenStudy (shayaan_mustafa):

no... did you use it correctly?

OpenStudy (anonymous):

i think so. what did i do wrong?

OpenStudy (shayaan_mustafa):

here your brackets are wrong. they should look like this. 3(3^2)t and 3(3)t^2

OpenStudy (shayaan_mustafa):

do you want to learn basic of binomial theorem?

OpenStudy (anonymous):

sure :D

OpenStudy (shayaan_mustafa):

ok. it will take much time. let us start. do you know the formula (a+b)^2=a^2+2ab+b^2

OpenStudy (anonymous):

yea

OpenStudy (shayaan_mustafa):

good. it is also a binomial theorem. Notice that here power two indicates that when we expand our formula, we get one more term than the no. of exponent. Here exponent is 2 while when we expanded then we got 3 terms.

OpenStudy (anonymous):

ok

OpenStudy (shayaan_mustafa):

similarly, I hope you know the formula (a+b)^3=a^3+3a^2b+3ab^2+b^3 here exponent is 3 while expanded terms are 4. This implies that expanded terms will always be one more than the exponent. Now is it clear?

OpenStudy (anonymous):

oo I think I get it now. Thank you.

OpenStudy (shayaan_mustafa):

welcome. ok now make a step ahead. here is binomial formula. \[(a+b)^{n}=C _{0}^{n}a ^{n}+C _{1}^{n}a ^{n-1}b+C _{2}^{n}a ^{n-2}b ^{2}.....\] Where C stands for combination. in general this combination is written as, \[C _{r}^{n}=n!/(r!(n-r)!)\] now in above formula is constantly increasing, r=0,r=1,r=2.......